# A semi-supervised framework for diverse multiple hypothesis testing scenarios

### v2

@@ 6.2 Real data experiments with p-values
We next evaluate RESET and related methods on a collection of publicly available datasets with p-values that were investigated by Zhang et al. [51] when introducing AdaFDR and by Lei and Fithian [33] when introducing AdaPT. Since AdaFDR and AdaPT operate under essentially the same assumptions as RESET’s, we expect these datasets to therefore satisfy Assumption 3 with the exception of two datasets, which we discuss at the end of this section. We provide a brief summary of each dataset below.
  • Three RNA-seq datasets (Airway [22], Bottomly [4], Pasilla [6]): The goal is to identify genes that are associated with varying gene expression levels in response to a change of conditions. In the Airway data, the differential expression analysis is conducted in response to a drug, dexamethasone; in the Bottomly data, the analysis is conducted between two mouse strains; and in the Pasilla data, the analysis is conducted between normal and so-called ‘Pasilla knockdown’ conditions [23]. In all cases the side information consists of the log-normalized gene counts.
  • A proteomics dataset (Proteomics [14]): The goal is to identify proteins that have different protein abundances when treated with rapamycin versus DMSO in yeast cells. Each protein consists of side information equal to the natural log of the number of peptides ‘quantified’ across all samples, i.e., the number of peptides belonging to the protein with their abundances measured.
  • Two microbiome datasets (Microbiome Enigma Ph and Microbiome Enigma Al [45, 31]): The goal is to identify microbiome organisms, recorded as OTUs (Operational Taxonomic Units), that are associated with certain conditions (pH and Al). Each hypothesis (OTU) is equipped with a two-dimensional side information — the ubiquity, which is defined as the percentage of samples detected with the OTU, and the mean nonzero abundance, which is defined as the average abundance of each OTU across the samples in which it was detected.
  • Two fMRI datasets (Auditory and Imagination) [48]: The goal is to identify voxels that are activated in response to different stimuli. In the first dataset, a single individual received auditory stimulus and in the second dataset, a single individual was instructed to imagine playing tennis. Each hypothesis (voxel) is equipped with four dimensional side information: the spatial position of the voxel (in three dimensions) and a categorical variable with the Brodmann label [5], which is used to delineate different areas of the brain with different functions.
  • Two eQTL datasets (Subcutaneous and Omentum [11, 10]): The goal is to identify single nucleotide polymorphisms (SNPs) that are associated with changes in gene expression levels. Such an association is referred to as an expression quantitative trait locus (eQTL). Each hypothesis (SNP) is equipped with the following four side information variables: (1) the distance between the SNP and the gene transcription’s start site (TSS), (2) the log gene expression level, (3) the alternative allele frequency (AAF) of the SNP and (4) the chromatin state of the SNP.
  • Gene-drug response data (Estrogen [12, 36]): The goal is to identify genes that are differentially expressed in response to a low-dosage estrogen treatment applied to breast cancer cells. Each hypothesis (gene) is assigned a ‘rank’ that was generated using an analysis of the gene expression levels in response to (1) a high-dosage estrogen treatment and (2) a medium-dosage estrogen treatment. Here, the ranks are in terms of the strength of association between the expression level of each gene and the level of dosage in (1) and (2). Thus, we can analyze the data twice, once for each side information variable based on (1) and (2).
Figure 12: Comparison of RESET and other methods using a collection of publicly available datasets. Each method was evaluated at the FDR thresholds of 1%, 5%, 10%, and 20%. The number of discoveries reported by each RESET method is averaged over 10 applications at FDR threshold and dataset. Each color and shape combination corresponds to a unique method. For readability, each point is jittered in the horizontal direction.
The results of each method applied to the collection of the above datasets is visually summarized in Figure 12. We can clearly see that RESET Ensemble is comparable to the other methods in terms of discoveries. Indeed, to argue this more accurately, we ranked each method across each dataset and at each FDR threshold considered (ties were averaged). Then we calculated the median rank for each method. RESET Ensemble and AdaPTg-GMM were tied with the highest median rank. Moreover, for each dataset and each FDR threshold in Figure 12, we computed the maximum number of discoveries across all methods and determined the relative power of each method to this maximum. The boxplots in Figure 13 display the relative powers for each method. Visually, RESET Ensemble and AdaPTg-GMM have the highest medians, with RESET Ensemble at 95.9% and AdaPTg-GMM at 94.5% with RESET Ensemble’s boxplot edging higher overall.
Figure 13: Relative power of each method. Each boxplot describes the relative power of each method computed for each dataset and FDR threshold in Figure 12.
@@
There are a couple of implementation details we would like to highlight here (for full details, see Supplementary Sections S.6-S.10). First, we used the same default options of RESET Ensemble on all datasets, except for the two fMRI datasets which exhibit a spike of p-values close to 1. In this case, it is sensible to define asymmetric target and decoy regions, $[0,0.3)$ and $(0.3,0.9]$ respectively as Chao and Fithian [8] did in their application of AdaPTg-GMM to this dataset. We also point out that we applied AdaPTg-GMM and ZAP using the p-value information rather than the test statistic information (see Section 3.1 for background on these methods). To do so, AdaPTg-GMM allows the user to directly input p-values instead of z-statistics. For ZAP, we converted the p-value information into z-values by defining $z_{i}=\pm\Phi^{-1}(p_{i}/2)$ of each p-value, $p_{i}$, where the signs are chosen i.i.d. uniformly. In other words, we input plausible z-values that correspond to the same p-value information used by the other tools. As we mentioned in Section 3, our implementation of RESET focuses on the use of p-values, and we leave the extension to test statistic information as part of future work.
@@
To evaluate the computation times for each method, we used the two fMRI datasets, which uses the largest number of side information variables and has a large number of hypotheses. Figure 14(A-B) shows the computation times for each considered method, where the fastest are the collection of RESET methods, ZAP and AdaFDR. RESET NN, RESET GAM and RESET RF are slightly faster than our ensemble approach which considers all of them.
@@
Figure 14: Comparison of runtime and power. In (A) and (B), we compare the computation time of each method on the fMRI datasets. We averaged the times of each RESET method over 10 applications. In (C) and (D), we compare the number of discoveries using the two eQTL datasets. The number of discoveries reported by each RESET method is averaged over 10 applications at FDR threshold and dataset. In all panels, we evaluated each method at FDR thresholds of: 1%, 5%, 10%, 20%.
Notably, we have excluded the analysis of the two eQTL studies from Figure 12. This is because this dataset exhibits a high degree of p-value dependency, even among the true null hypotheses and thus violates Assumption 3 which guarantees our type-1 error control. Indeed, this phenomenon is due to linkage disequilibrium in which SNPs behave dependently. As a consequence, the dataset exhibits virtually identical copies of the same p-value, with similar side information, even if the p-value is large, suggesting a true null hypothesis. Figure 14(C-D) shows that RESET Ensemble reports substantially more discoveries on these eQTL datasets than all the non-RESET variants. This is because many estimating and training decoys are virtually the same and the random forest is sensitive to this leakage of information. Evidently, RESET NN is more robust to this problem: the number of discoveries is in the same ‘ballpark’ as other methods. Note that RESET GAM shares identical results with RESET RF since it fails to perform a spline basis expansion on one of the side information variables (the chromatin state of the SNP), and so RESET GAM switches to RESET RF in this case.
6.3 FDP control with p-values
@@
Figure 15: Comparison of RESET and other methods with FDP control using a collection of publicly available data. Each panel plots the number of discoveries obtained by RESET Ensemble, FDP-SD and GR-SD at a range of FDR thresholds (1%, 5%, 10%, 20%) with varying confidence $1-\gamma$ (0.5, 0.8, 0.9).
@@
In this final section, we demonstrate RESET’s application of FDP control in the p-value setting. Here we define the following asymmetric target and decoy regions, $[0,0.3)$ and $(0.3,0.9]$ respectively, for all datasets. The rationale stems from our later analysis in the Discussion, as well as our comments at the end of Section 5.2.3, which point out that when the ratio of estimating decoys to true null targets is small, the subsequent FDP-SD step in RESET might incur a loss of power. Hence, to rectify this, we essentially double the number of estimating decoys under consideration by defining a decoy region that is twice as large as the corresponding target region.
As far as we can tell, there are no p-value based testing procedures that use side information while controlling the FDP in the sense of Section 2. Hence, we compare RESET to the following generic methods for FDP control. The first of these methods is the previously considered FDP-SD that we analyzed in the competition setting, but adapted for the p-value setup. Each of the labels and scores $(L,W)$ required for FDP-SD are constructed in the same way as in RESET’s Step 1 (Section 4), where we used the same asymmetric target and decoy regions described above. The second method we compare to is the p-value based analogue called GR-SD [20] (see Section 3).
@@
Figure 15 displays the number of discoveries obtained by each method on the collection of real datasets we considered in the previous section, with varying FDR thresholds (1%, 5%, 10%, 20%) and confidence levels (50%, 80%, 90%). RESET Ensemble is much more powerful than the alternative non-side information approaches with a few exceptions.
@@ 7 Discussion
We introduced RESET, a flexible wrapper for any semi-supervised learning algorithm. Alongside, we provide an implementation of RESET, called RESET Ensemble, that is power-wise competitive with recently developed methods for competition as well as p-value based multiple testing problems with side information, fast compared to most tools, and flexible as it allows finite-sample of either FDR or FDP control.
@@
In terms of power, we highlight the fact that competing tools, such as the variants of AdaPT or Adaptive Knockoffs, require an initial selection of one of those variants. Making a selection after observing the discovery lists of each tool will generally compromise its FDR control. Thus, we compared RESET Ensemble to the default competing method if one was available, which was the case of Adaptive Knockoff’s GAM filter, or separately to each variant, which was the case of AdaPT. Based on our comparisons, we conclude that RESET Ensemble is able to achieve comparable or slightly greater power than many of these tools while freeing the user from having to choose the ‘right’ method to begin with.
@@
As far as computational runtime, we found that our collection of RESET methods provide a faster alternative to all competing methods with finite FDR control in our considered experiments. The only case in which RESET was surpassed in terms of speed was ZAP-asymp, which demonstrated marginally faster runtimes than RESET Ensemble in one of the fMRI datasets (Figure 14B). However, ZAP-asymp’s fast runtimes come at the cost of asymptotic FDR control. The most striking difference in runtimes was in the peptide detection context using mass spectrometry data, in which all variants of Adaptive Knockoffs were generally too slow to be practical.
@@
Finally, we point out that RESET’s compatibility with FDR or FDP control is unique. In the competition setting, we note that it would be challenging to combine Adaptive Knockoffs with FDP-SD because FDP-SD is a stepdown procedure while Adaptive Knockoffs is essentially a step-up procedure. Katsevich and Ramdas [27] provide a general method for adaptive ordering using their method for simultaneous FDP bounds. However, (a) they do not provide an exact implementation of this, only a theoretical discussion, and (b) their approach is conservative since it controls the FDP at every cutoff (which allows the analyst to explore the nested discovery lists without breaking FDP control). To the best of our knowledge, in the competition framework with side information, RESET is the only method of reordering that may be flexibly used with FDP-SD without the additional cost of (b). Similarly, we know of no method that allows finite sample FDP control in the p-value setting with side information.
On the topic of RESET’s FDP control, we point out the following limitation. When the number of discoveries is small, e.g., when the confidence is too high, RESET may yield less power, even if the side information is reasonably informative. To illustrate this phenomenon, we computed the bounds, $\delta_{i}$ on the number of decoy wins in the top $i$ scores from Algorithm S2, which are used to determine the reported list of discoveries in FDP-SD (with $c=1/2$) and RESET (using the default $c=2/3$). We considered a confidence level of $1-\gamma=90\%$ and an FDP threshold of $\alpha=1\%$. Since we removed approximately half the decoys for training purposes in RESET, the bounds need to be doubled to fairly compare between the two approaches.
Refer to caption Refer to caption
Figure 16: RESET’s and FDP-SD’s bounds on the number of decoy wins. We recorded, on the left y-axis, the bounds used to determine the list of FDP-SD discoveries and twice the bounds used to determine RESET’s discoveries at a range of indices at $\alpha=1\%$ and $1-\gamma=90\%$. On the right y-axis, we plot the ratio of these bounds (in black). The left panel looks at indices between 501 to 2K, while the right panel looks at indices from 10K to 20K. An index corresponds to the number of top scoring hypotheses (regardless of their labels).
Figure 16 shows the bounds for FDP-SD, and double the bounds for RESET, along with the ratio of the two at a range of different indices. We can see that for smaller indices, this ratio is much smaller, initially at 0% at an index of 501 and 50% at an index of 1K, while at 10K this ratio is much higher at 94%. If the order of the hypotheses is the same, then RESET’s smaller bounds imply it will report fewer discoveries because RESET employs FDP-SD to control the FDP by searching for the largest index $i$ s.t. $D_{i_{0}}\leq\delta_{i},\dots,D_{i}\leq\delta_{i}$, and reports the top $i$ scoring positively labeled hypotheses. In other words, if RESET is to report roughly the same number of discoveries as FDP-SD, then RESET needs to rearrange the targets and estimating decoys successfully enough so as to make up for a 50% difference in the bounds when considering the top 1K scores, while only 6% when considering the top 10K. If the side information is not rich enough, then it is possible RESET will report few discoveries than FDP-SD in these cases. We intend to investigate this as part of future work.
Figure 17: Variability in the number of discoveries. in the Airway data and in Simulation 5 (A) Histogram of the number of RESET Ensemble discoveries applied 100 times to the Airway data with a 10% FDR threshold and varying its internal seed. (B) Boxplots of RESET’s and AdaPT’s power in 100 applications of the ‘circle in the middle’ setup of Simulation 5 using a 10% FDR threshold. There are two outliers in RESET’s case (which occurred when the signal was considerably low).
@@
Notably, RESET’s discovery list is variable since it randomly splits the decoys into two sets: one set for training and the other for estimating the false discoveries. This variability is demonstrated in the histogram of the number of discoveries in Figure 17A using 100 applications of RESET Ensemble to the Airway dataset at an FDR threshold of 10%. The resulting number of discoveries has a mean of 5998 with a standard deviation of about 85. While this is problematic to an extent, keep in mind that the data distribution is random, and the added variance from RESET may be marginal at times. Indeed, in Figure 17(B), we find that in the case of Simulation 5 from Section 6.1.1, RESET is overall arguably less variable than AdaPT, even though AdaPT does not employ any internal randomization. In practice, users may get around RESET’s randomization by setting the random number generator to be a function of the input data and parameters. We plan on adding this feature to a future update.
Lastly, while RESET is a flexible method for the competition and p-value based testing, we wish to further extend it to the test-statistic setting. There is more than one way of generalizing RESET to accommodate a test statistic $z_{i}$ for each hypothesis $H_{i}$. Moreover, there are several ‘types’ of interactions between the side information $\mathbf{x}_{i}$ and the test statistic $z_{i}$ which are not preserved when switching to p-values, as outlined by Leung and Sun [35]. Thus it is not clear that RESET Ensemble will be equally applicable in this setting without substantial modifications.
@@
S.1 Algorithms
Algorithm S1 Selective SeqStep / SeqStep+ (adopted from Selective Sequential Step+ of [1])
@@
$\bullet$ $c\in(0,1)$ the probability of a null target/feature win;
$\bullet$ $\alpha\in(0,1)$ the FDR threshold;
@@
3:sort the paired $(L_{i},W_{i})$ in decreasing order of $W_{i}$triangleright ties are randomly broken
4:$D_{k}\leftarrow\#\{i\leq k:L_{i}=-1\}$triangleright number of decoy/knockoffs wins in top $k$ scores
5:$T_{k}\leftarrow\#\{i\leq k:L_{i}=1\}$triangleright number of target/original feature wins in top $k$ scores
@@
7:     $k_{0}\leftarrow\max\{k:\frac{D_{k}}{T_{k}\vee 1}\cdot\frac{c}{1-c}\leq\alpha\}$
@@
9:     $k_{0}\leftarrow\max\{k:\frac{D_{k}+1}{T_{k}\vee 1}\cdot\frac{c}{1-c}\leq\alpha\}$
@@
11:return $R_{\alpha}\leftarrow\{i:\text{the (pre-sorted) $W_{i}$ is in the top $k_{0}$ ranks and $L_{i}=1$}\}$
Algorithm S2 FDP-SD (adopted from [40])
@@
$\bullet$ $c\in(0,1)$ the probability of a null target/feature win;
$\bullet$ $\alpha\in(0,1)$ the FDP threshold;
$\bullet$ $\gamma\in(0,1)$ for a $1-\gamma$ confidence level;
@@
3:sort the paired $(L_{i},W_{i})$ in decreasing order of $W_{i}$triangleright ties are randomly broken
4:$\lambda\leftarrow 1-c$triangleright the probability of a null decoy win
5:$R\leftarrow(1-\lambda)/(c+1-\lambda)$
6:$i_{0}\leftarrow\max\{1,\lceil(\lceil\log_{1-R}(\gamma)\rceil)/\alpha\rceil\}$
@@
8:     $D_{i}\leftarrow\#\{j\leq i:L_{j}=-1\}$triangleright number of decoy wins in top $i$ scores
9:     $\delta_{i}\leftarrow\max\{d\in\{0,1,\dots,i\}:F_{B(\lfloor(i-d)\alpha+1+d,R\rfloor)}(d)\leq\gamma\}$triangleright $F_{B(n,p)}$ denotes the CDF of a Binomial$(n,p)$ RV
@@
39:$R_{\alpha,\gamma}\leftarrow\{i:\text{the (pre-sorted) $W_{i}$ is in the top $k_{FDP}$ ranks and $L_{i}=1$}\}$
Algorithm S3 GR-SD (adopted from [20])
@@
$\bullet$ $\alpha\in(0,1)$ the FDP threshold;
$\bullet$ $\gamma\in(0,1)$ for a $1-\gamma$ confidence level;
@@
6:     $\delta_{i}\leftarrow F^{-1}[k_{i},m-i+1](\gamma)$triangleright $F^{-1}[\alpha,\beta](\cdot)$ denotes the CDF of a Beta$(\alpha,\beta)$ RV
@@
8:$k_{GR}\leftarrow\max\{i:\prod_{j=1}^{i}1_{p_{i}\leq\delta_{i}}=1\text{ or }i=0\}$
9:$R_{\alpha,\gamma}\leftarrow\{i\in[m]:i\leq k_{GR}\}$
Algorithm S4 RESET
@@
$\bullet$ $\{(L_{i},W_{i},\mathbf{x}_{i})\,:\,i=1,\dots,m\}$ each hypothesis’ (label, winning score, side information);
$\bullet$ $\alpha$ FDR threshold for the discovery list;
$\bullet$ $s$ the probability of assigning a decoy to the training set (default: $s=1/2$);
$\bullet$ $f$ a semi-supervised machine learning model;
$\bullet$ $c_{0}$ an upper bound probability for a true null label $\mathbb{P}(L_{i}=1)\leq c_{0}$ ($i$ is a true null);
$\bullet$ isFDR - a boolean which determines whether FDR or FDP control is desired;
$\bullet$ $\gamma$ a confidence parameter if FDP control is desired;
@@
3:$I\leftarrow$ a subset of the decoy win indices, $\{i:L_{i}=-1\}$, determined by randomly including each one with probability $s$
4:$\tilde{L}_{i}\leftarrow-1$ for $i\in I$triangleright training decoys
5:$\tilde{L}_{i}\leftarrow 1$ for $i\in J:=I^{c}$triangleright pseudo-targets
6:$(\tilde{W}_{i})_{i=1}^{m}\leftarrow f\left((W_{i},U_{i},\tilde{L}_{i})_{i=1}^{m}\right)$triangleright where $f$ is any machine learning model
@@
8:     $R\leftarrow\text{SSS+}((\tilde{W}_{i},L_{i})_{i\in J},c=\frac{c_{0}}{1-s\cdot(1-c_{0})},\alpha)$triangleright Selective SeqStep+
@@
10:     $R\leftarrow\text{FDP-SD}((\tilde{W}_{i},L_{i})_{i\in J},c=\frac{c_{0}}{1-s\cdot(1-c_{0})},\alpha,\gamma)$triangleright FDP-stepdown

### v3

@@ 6.4 Real data experiments with p-values
We next evaluate RESET and related methods on a collection of publicly available datasets with p-values. These were investigated by Zhang et al. (2019), when introducing AdaFDR, and by Lei and Fithian (2018) when introducing AdaPT. A description of each dataset can be found in Section D.3.
As we saw above, RESET Ensemble and the other tools we analyze here exhibit some robustness to deviations from the independence assumption that is built into Assumptions 3 and 4. However, despite this fact, we sought to remove the two Adipose datasets from this analysis because RESET’s heuristics, described in Section C.2.4, flagged these datasets for high dependency (cf. Section F.3).
Figure 6: Power comparison. Each boxplot is made of 40 data points showing the relative power of a method on all combinations of dataset and FDR threshold. There are ten dataset as described in Section D.3 (excluding the two Adipose datasets) and four FDR thresholds: 1%, 5%, 10% and 20%. RESET Ensemble’s power and its variants were averaged over 10 applications.
For each dataset (excluding the two Adipose datasets) and each FDR threshold in $\{1\%,5\%,10\%,20\%\}$, we computed the maximum number of discoveries across all methods and determined the power of each method relative to this maximum (Figure 6). RESET Ensemble (94.4%), RESET NN (93.6%) and AdaPTg-GMM (91.5%) have the highest medians, with RESET Ensemble also arguably having the tightest boxplot. Figure E6 shows the number of discoveries for each method on each of the 40 combinations of dataset and FDR threshold.
@@
There are a couple of implementation comments we highlight here (see Sections C.7-C.11 for details). First, we used the same default options of RESET Ensemble for all datasets, except for the two fMRI datasets which exhibit a spike of p-values close to 1. To address this, we followed Chao and Fithian (2021)’s suggestion of defining asymmetric target and decoy regions of $[0,0.3)$ and $(0.3,0.9]$ respectively. Second, we applied AdaPTg-GMM and ZAP using the p-value information rather than the test statistic information. To do so, AdaPTg-GMM allows the user to directly input p-values instead of z-statistics. For ZAP, we converted the p-value information into z-values by defining $z_{i}=\pm\Phi^{-1}(p_{i}/2)$ for each p-value, $p_{i}$, where the signs are chosen i.i.d. uniformly. In other words, we input plausible z-values that correspond to the same p-value information used by the other tools. As we mentioned in Section 3, our implementation of RESET focuses on the use of p-values, and we leave the extension to test statistic information as part of future work.
@@
To evaluate the runtime of each method, we used the two fMRI datasets, which use the largest number of side information variables and has a large number of hypotheses. Figure 7(A-B) shows the runtimes for each considered method, where the fastest are the collection of RESET methods, ZAP and AdaFDR. RESET NN, RESET GAM and RESET RF are slightly faster than our ensemble approach which considers all of them.
@@
Figure 7: Runtime comparison. In (A) and (B), we compare the runtime of each method on the fMRI datasets. We averaged the times of each RESET method over 10 applications.
6.5 FDX control with p-values
@@
Figure 8: Comparison of RESET and other methods with FDX control using a collection of publicly available data. Each panel shows the number of discoveries obtained by RESET Ensemble, FDP-SD and GR-SD at a range of FDR thresholds (1%, 5%, 10%, 20%) with varying confidence $1-\gamma$ (0.5, 0.8, 0.9). RESET Ensemble and FDP-SD were averaged over 10 applications.
@@
In this final section, we demonstrate RESET’s FDX control in the p-value setting. To the best of our knowledge, there are no existing p-value-based testing procedures that offer finite sample FDX control while utilizing side information. Therefore, we compare RESET Ensemble against FDP-SD (Algorithm A2) and GR-SD (Algorithm A3). The labels and scores $(L,W)$ used in RESET Ensemble and FDP-SD are defined using the asymmetric target and decoy regions—$[0,0.3)$ and $(0.3,0.9]$, respectively—for all datasets. The rationale stems from our analysis in Section F.4, which points out that when the ratio of estimating decoys to true null targets is small, RESET might incur a loss of power when controlling the FDX. To rectify this, we define a decoy region that is twice as large as the corresponding target region.
@@
Figure 8 displays the number of discoveries obtained by each method on the collection of real datasets we considered in the previous section. The numbers are averaged over 10 runs while varying the FDR thresholds (1%, 5%, 10%, 20%) and the confidence levels (50%, 80%, 90%). With a few exceptions RESET Ensemble is generally much more powerful than the alternative approaches that make no use of the side information.
@@ 7 Discussion
RESET provides a general framework for FDR and FDX control through a flexible wrapper for any semi-supervised learning algorithm. Alongside, we introduced an implementation of RESET, called RESET Ensemble, that is power-wise competitive with recently developed methods for competition as well as p-value based multiple testing problems with side information, is fast compared to most tools, and is flexible allowing finite sample FDR and FDX control.
@@
In terms of power, we highlight the fact that competing tools, such as the variants of AdaPT or Adaptive Knockoffs, require an initial selection of one of those variants. Making a selection after observing the discovery lists of each tool will generally compromise its FDR control. Thus, we compared RESET Ensemble to the default competing method if one was available, which is the case with Adaptive Knockoff’s GAM filter, or separately to each variant, which is the case for AdaPT. Based on our comparisons, we conclude that RESET Ensemble is able to achieve comparable or slightly greater power than many of these tools while freeing the user from having to choose the ‘right’ method to begin with.
@@
We found that each of our RESET methods was consistently faster than any of the competing methods for finite FDR control with side-information we considered here. More generally, the only case in which RESET was surpassed in terms of speed was ZAP-asymp, which demonstrated marginally faster runtimes than RESET Ensemble in one of the fMRI datasets (Figure 7B). However, ZAP-asymp’s fast runtimes come at the cost of asymptotic FDR control. The most striking difference in runtimes was in the peptide detection context using mass spectrometry data, in which all variants of Adaptive Knockoffs were generally too slow to be practical. RESET’s impressive speed compared to Adaptive Knockoffs and the AdaPT variants is due to the fact that these methods work by iterating between training a classification algorithm, and deciding which hypothesis to remove from the candidate set $\mathcal{\hat{S}}_{t}$ as explained in Section 2.1. In contrast, RESET Ensemble can be heavily parallelized, and only performs two iterations.
@@
Finally, we point out that RESET’s compatibility with FDR as well as FDX control is unique. In the competition setting, we note that it would be challenging to combine Adaptive Knockoffs with FDP-SD because FDP-SD is a stepdown procedure while Adaptive Knockoffs is essentially a step-up procedure. Katsevich and Ramdas (2020) provide a general method for adaptive ordering using their method for simultaneous FDP bounds. However, (a) they do not provide an exact implementation of this, only a theoretical discussion, and (b) their approach is conservative since it controls the FDX at every cutoff (which allows the analyst to explore the nested discovery lists without breaking FDX control). To the best of our knowledge, in the competition framework with side information, RESET is the only method of reordering that may be flexibly used with FDP-SD without the additional cost of (b). Similarly, we know of no method that allows finite sample FDX control in the p-value setting with side information.
@@
Notably, RESET’s discovery list is variable since it randomly splits the decoys into two sets: one set for training and the other for estimating the false discoveries. While this is problematic to an extent, keep in mind that the data distribution is random, and the added variance from RESET may be marginal at times (cf. Section F.5).
Lastly, we point out a few avenues for further research. While RESET is a flexible method for the competition and p-value based testing, we wish to further extend it to the test-statistic setting. There is more than one way of generalizing RESET to accommodate a test statistic $z_{i}$ for each hypothesis $H_{i}$. Moreover, there are several ‘types’ of interactions between the side information $\mathbf{x}_{i}$ and the test statistic $z_{i}$ which are not preserved when switching to p-values, as outlined by Leung and Sun (2022). Thus it is not clear that RESET Ensemble will be equally applicable in this setting without substantial modifications. Second, RESET relies on a handful of prespecified quantities: the probability $s$ of assigning decoys to the estimation set, and the definitions of the target and decoy regions as $[0,a)$ and $(b_{1},b_{2}]$, respectively. Moreover, in Step (1) we employ a mirroring function that converts larger $p$-values in the decoy region into smaller values in the target region. Given that the true null p-values are possibly non-uniform, we could consider other masking functions as suggested by Chao and Fithian (2021). The selection of these parameters and the masking function would ideally be data-driven which we leave as part of future work. Finally, we may consider alternatives to our semi-supervised learning engine, RESET Ensemble. One direction is to pair RESET’s data-splitting scheme with AdaPT’s iterative fitting procedure, which could further increase power (at the expense of additional computation time). Another is to broaden RESET Ensemble’s library of classifiers to include the two-group models that are used by many of the competing methods.
@@
Appendix A Algorithms
Algorithm A1 Selective SeqStep / SeqStep+ (adopted from Selective Sequential Step+ of Barber and Candès (2015))
@@ Appendix A Algorithms
$\bullet$ $c\in(0,1)$ - the probability of a null target/feature win;
$\bullet$ $\alpha\in(0,1)$ - the FDR threshold;
@@
3:sort the paired $(L_{i},W_{i})$ in decreasing order of $W_{i}$ $\triangleright$ ties are randomly broken
[…]
@@
7:  $\tau\leftarrow\max\{t:\frac{A_{t}}{R_{t}\vee 1}\cdot\frac{c}{1-c}\leq\alpha\}$
@@
9:  $\tau\leftarrow\max\{t:\frac{A_{t}+1}{R_{t}\vee 1}\cdot\frac{c}{1-c}\leq\alpha\}$
@@
[…]
Algorithm A2 FDP-SD (adopted from Luo et al. (2023))
@@
$\bullet$ $c\in(0,1)$ - the probability of a null target/feature win;
$\bullet$ $\alpha\in(0,1)$ - the FDP threshold;
$\bullet$ $\gamma\in(0,1)$ - for a $1-\gamma$ confidence level;
@@
3:sort the paired $(L_{i},W_{i})$ in decreasing order of $W_{i}$ $\triangleright$ ties are randomly broken
[…]
@@
8:  $A_{i}\leftarrow\#\{j\leq i:L_{j}=-1\}$ $\triangleright$ number of decoy wins in top $i$ scores
9:  $\delta_{i}\leftarrow\max\{d\in\{0,1,\dots,i\}:F_{B(\lfloor(i-d)\alpha+1+d,R\rfloor)}(d)\leq\gamma\}$ $\triangleright$ $F_{B(n,p)}$ denotes the CDF of a Binomial(n,p)(n,p) RV
@@
[…]
Algorithm A3 GR-SD (adopted from Guo and Romano (2007))
@@
$\bullet$ $\alpha\in(0,1)$ - the FDP threshold;
$\bullet$ $\gamma\in(0,1)$ - for a $1-\gamma$ confidence level;
@@
[…]
@@
[…]
Algorithm A4 RESET
@@
$\bullet$ $\{(L_{i},W_{i},\mathbf{x}_{i})\,:\,i=1,\dots,m\}$ - each hypothesis’ (label, winning score, side information);
$\bullet$ $\alpha$ - FDR threshold for the discovery list;
$\bullet$ $s$ - the probability of assigning a decoy to the training set (default: $s=1/2$);
$\bullet$ $f$ - a semi-supervised machine learning model;
$\bullet$ $c_{0}$ - an upper bound probability for a true null label $\mathbb{P}(L_{i}=1)\leq c_{0}$ (ii is a true null);
$\bullet$ isFDR - a boolean which determines whether FDR or FDX control is desired;
$\bullet$ $\gamma$ - a confidence parameter if FDX control is desired;
@@
[…]
@@
8:  $R\leftarrow\text{SSS+}((\widetilde{W}_{i},L_{i})_{i\in J},c=c_{e}:=\frac{c_{0}}{1-s\cdot(1-c_{0})},\alpha)$ $\triangleright$ Selective SeqStep+
@@
10:  $R\leftarrow\text{FDP-SD}((\widetilde{W}_{i},L_{i})_{i\in J},c=c_{e}=\frac{c_{0}}{1-s\cdot(1-c_{0})},\alpha,\gamma)$ $\triangleright$ FDP-stepdown